Two investments were made totaling $3,000. For a certain year, these investments yielded $216 in simple interest. Part of the $3,000 was invested at 6% and part at 8%. Find the amount invested at each rate.
Answer by rfer(16322)

.06x+.08(3000-x)=216
.06x+240-.08x=216
-.02x=-24
x=$1200 at 6%
3000-x=$1800 at 8%